Operation Summary for June 4, 2010 - June 7, 2010 |
The Integrated Luminosity for the period from 5/31/10 to 6/7/10 was 68.54 pb-1. Friday June 4 The day shift began with Operations monitoring the Recycler stash (365.3E10) and the Antiproton Source (Pbar) stack (23.0E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store. During the day shift, Operations put the TeV into shot setup at 8:12 AM and established store 7870 at 9:15 AM with an average initial luminosity (AIL) of 360.1E30. Operations reported that NuMI was off from 1:40 PM to 2:35 PM to investigate a high temperature alarm on a LCW system. Operations reported other short downtimes due to Booster ion pump problems. During the evening shift, Operations established beam to the FTBF MINERvA experiment T-977 for about one hour starting at 5 PM. Operations terminated store 7870 at 10:49 PM and put the TeV into shot setup. Operations reported that a TeV RF station (TRF2) faulted for a couple of seconds, but did not trip off. CDF had to turn off for a few minutes due to beam in the abort gap. Operations reported several downtimes due to the following: Booster ion pump (IPRF14) trips and to Linac RF station (LRF5) trips. Saturday June 5 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (6.6E10), and the stack (46.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 7871 at 12:12 AM with an AIL of 356.61E30. Operators search and secured the MTest section #2 area. IPRF14 tripped off several more times. Operations reported that FESS turned off the power to Wilson hall as scheduled at 6 AM. Operations established beam to T-977 at 7:09 AM. During the day shift, Operations terminated store 7871 at 2 PM, put the TeV into shot setup, and established store 7872 at 3:20 PM with an AIL of 349E30. MiniBooNE reported having a network problem. During the evening shift, FESS turned on the Wilson Hall power at 5:30 PM. IPRF14 caused two Booster RF stations (BRF1 and 2) to trip off twice in three hours. Sunday June 6 The midnight shift began with Operations monitoring store 7872 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Controls personnel worked on the MiniBooNE network and data issues. IPRF14 tripped off several more times. Operations terminated store 7872 at 4:41 AM, put the TeV into shot setup, and established store 7873 at 6:19 AM with an AIL of 344.72E30. Operations established beam to T-977 at 7:41 AM. During the day shift, operators tuned the Proton Source. Cryo system technicians worked on the TeV sector C3 wet engine flywheel. During the evening shift, Operations terminated store 7873 at 8:52 PM, put the TeV into shot setup, and established store 7874 at 10:13 PM with and AIL of 354.18E30. Monday June 7 The midnight shift began with Operations monitoring store 7874 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported a differential pressure alarm for the F1 LCW system; Mechanical Support personnel came in and valved in the back up filter. Operations accessed Switchyard enclosures F1 and G2 to investigate high water alarms; Operators breakered on the F1 sump pump and FESS personnel replaced a fuse for the G2 sump pump. Operations established beam to T-977 at 4:24 AM. IPRF14 continued to trip off and be reset. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 06/07/2010 email Fermilab |
| Security, Privacy, Legal | |